# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ A Dirichlet form $(\mathscr{E},\mathscr{D})$ on $L^{2}(\nu)$ is called quasi-regular if:
13.3. Quasi-regularity of $m$-labeled Dirichlet forms.
Let $\mu$ be a random point field on $\mathbb{R}^{d}$. Let $\mu^{[m]}$ be the $m$-reduced Campbell measure. Let
$\mathscr{E}^{\mu^{[m]}}(f,g)=\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}\mathbb{D}_{\mathfrak{a}}^{[m]}[f,g]d\mu^{[m]},$
where $\mathfrak{a}$ satisfies (2.39) and $\mathbb{D}_{\mathfrak{a}}^{[m]}$ is defined by (7.5). Let $\mathscr{D}_{\circ}^{\mu^{[m]}}$ be as in (7.12):
$\mathscr{D}_{\circ}^{\mu^{[m]}}=\{f\in\mathscr{D}_{\circ}^{[m]};\mathscr{E}^{\mu^{[m]}}(f,f)<\infty,\,f\in L^{2}(\mu^{[m]})\}.$
We make assumptions:
A.1 $(\mathscr{E}^{\mu^{[m]}},\mathscr{D}_{\circ}^{\mu^{[m]}})$ is closable on $L^{2}(\mu^{[m]})$.
A.2 $\sum_{k=0}^{\infty}k\mu(S_{R}^{k})<\infty$
A.3 $\mu$ has a bounded $k$-density function on $\overline{S}_{R}^{k}$ for each $R,k\in\mathbb{N}$.
Let $(\mathscr{E}^{\mu^{[m]}},\overline{\mathscr{D}}^{\mu^{[m]}}_{\circ})$ be the closure of $(\mathscr{E}^{\mu^{[m]}},\mathscr{D}_{\circ}^{\mu^{[m]}})$ on $L^{2}(\mu^{[m]})$. We write $(\mathscr{E}^{\mu},\overline{\mathscr{D}}_{\circ}^{\mu})$ for $m=0$.
Theorem 13.1 ([22, Th. 1]).
Assume A.1 for $m=0$, A.2, and A.3. Then, the Dirichlet form $(\mathscr{E}^{\mu},\overline{\mathscr{D}}_{\circ}^{\mu})$ on $L^{2}(\mu)$ is quasi-regular.
@@ A.2 $\sum_{k=0}^{\infty}k\mu(S_{R}^{k})<\infty$
Theorem 13.1 was proved by [22, Th. 1]. In [22, Th. 1], $\mathfrak{a}$ was supposed to be the unit matrix. It is obvious that Theorem 13.1 follows from this case as well because, from (2.39), $\mathfrak{a}$ is uniformly elliptic and bounded. ∎
In the following theorem, we prove the quasi-regularity without assumption A.3.
@@
Assume A.1 for $m=0$ and A.2. Then, the Dirichlet form $(\mathscr{E}^{\mu},\overline{\mathscr{D}}_{\circ}^{\mu})$ on $L^{2}(\mu)$ is quasi-regular.
@@
((QR1)) and ((QR3)) follow from A.1 and A.2 in the exactly same fashion as the proof of [22, Th. 1].
In the proof of [22, Th. 1], A.3 was used only in the proof of ((QR2)). More specifically, the use of A.3 was limited to the proof of 2.2 found in [22, Lem. 2.4]. Here, 2.2 [22, Lem. 2.4] is the statement such that, for a $\sigma[\pi_{S_{R-\delta}}]$-measurable bounded function $f$, $0<\delta<R$,
$\lim_{\varepsilon\to 0}\lVert\mathcal{J}_{R,\varepsilon}f(\mathsf{s})-f\rVert_{L^{2}(\mu)}=0,$ (13.4)
where $\mathcal{J}_{R,\varepsilon}$ is defined by (13.6). Additionally, (13.4) was used solely to establish Lemma 2.4 3 in the same work, which asserts that
$\text{$\mathcal{D}_{\circ}^{\mu}$ is dense in $L^{2}(\mu)$}.$ (13.5)
Therefore, our goal is to demonstrate (13.5) without using A.3. To accomplish this, we will recall the mollifier $\mathcal{J}_{R,\varepsilon}$ on $\mathsf{S}$ that was introduced in [22].
Let $j\!:\!\mathbb{R}^{d}\!\to\!\mathbb{R}$ be a non-negative, smooth function such that $\int_{\mathbb{R}^{d}}j(x)dx=1$, $j(x)=0$ for $|x|\geq 1/2$. Let $j_{\varepsilon}(x)=\varepsilon^{d}j(x/\varepsilon)$ and $j_{\varepsilon}^{m}((x_{1},\ldots,x_{i}))=\prod_{j=1}^{m}j_{\varepsilon}(x_{j})$.
For a $\sigma[\pi_{S_{R}}]$-measurable function $f$, we set
$\mathcal{J}_{R,\varepsilon}f(\mathsf{s})=\begin{cases}j_{\varepsilon}^{m}*f_{R}^{m}(\mathbf{x}_{R}^{m}(\mathsf{s}))&\text{ for }\mathsf{s}\in\mathsf{S}_{R}^{m}\quad(m\geq 1),\\ f(\mathsf{s})&\text{ for }\mathsf{s}\in\mathsf{S}_{R}^{0}.\end{cases}$ (13.6)
Here, $f_{R}^{m}$ is $S_{R}^{m}$-representation of $f$ and $\mathbf{x}_{R}^{m}(\mathsf{s})$ is the $S_{R}^{m}$-coordinate introduced in Section 1. Additionally, * denotes the convolution in $(\mathbb{R}^{d})^{m}$, that is,
$j_{\varepsilon}^{m}*f_{R}^{m}(\mathbf{x})=\int_{(\mathbb{R}^{d})^{m}}j_{\varepsilon}^{m}(\mathbf{x}-\mathbf{y})f_{R}^{m}(\mathbf{y})d\mathbf{y},$ (13.7)
where we set $f_{R}^{m}(\mathbf{x})=0$ for $\mathbf{x}\not\in S_{R}^{m}$ and we assume $f_{R}^{m}\in L^{1}(S_{R}^{m},d\mathbf{x})$ to make (13.7) meaningful. We now prove that $\mathcal{D}_{\circ}^{\mu}$ is dense in $L^{2}(\mu)$.
Let $0<\delta<R$ ($\delta\in\mathbb{R}$) and $f\in C_{b}(\mathsf{S})\cap\sigma[\pi_{S_{R-\delta}}]$. Then $f$ is a bounded continuous, $\sigma[\pi_{S_{R-\delta}}]$-measurable function on $\mathsf{S}$ by definition. From [22, Lem. 2.4 2.1] we see $\mathcal{J}_{R,\varepsilon}f\in\mathcal{D}_{\circ}^{\mu}$ for $0<\varepsilon<\delta$. Moreover, because $f\in C_{b}(\mathsf{S})$, we see from (13.6) that
$\lim_{\varepsilon\to 0}\mathcal{J}_{R,\varepsilon}f(\mathsf{s})=f(\mathsf{s})\quad\text{ for each }\mathsf{s},$ (13.8)
$\sup_{\mathsf{s}\in\mathsf{S}}|\mathcal{J}_{R,\varepsilon}f(\mathsf{s})|\leq\sup_{\mathsf{s}\in\mathsf{S}}|f(\mathsf{s})|<\infty.$ (13.9)
From (13.8)–(13.9) we can apply the Lebesgue convergence theorem to obtain
$\lim_{\varepsilon\to 0}\int_{\mathsf{S}}|\mathcal{J}_{R,\varepsilon}f(\mathsf{s})-f(\mathsf{s})|^{2}\mu(d\mathsf{s})=\int_{\mathsf{S}}\lim_{\varepsilon\to 0}|\mathcal{J}_{R,\varepsilon}f(\mathsf{s})-f(\mathsf{s})|^{2}\mu(d\mathsf{s})=0$ (13.10)
for each $R\in\mathbb{N}$. Because $\cup_{0<\delta<R,\,\delta\in\mathbb{R},R\in\mathbb{N}}C_{b}(\mathsf{S})\cap\sigma[\pi_{S_{R-\delta}}]$ is dense in $L^{2}(\mu)$, we see that $\mathcal{D}_{\circ}^{\mu}$ is dense in $L^{2}(\mu)$ from (13.10). Thus, we have obtained (13.5) and completed the proof of Theorem 13.2. ∎
In Theorem 13.2 and Theorem 13.3, we significantly relax the assumptions introduced in Theorem 13.1. Specifically, if $(\mathscr{E}^{\mu^{[m]}},\mathscr{D}_{\circ}^{\mu^{[m]}})$ is found to be closable on $L^{2}(\mu^{[m]})$, then the existence of a density function is no longer necessary. The random point field $\mu_{\mathrm{GAF}}$, defined by the zeros of the planar Gaussian analytic function, does not possess a density function due to its number and mean rigidities [8, 9]. Therefore, the stochastic dynamics that are reversible with respect to $\mu_{\mathrm{GAF}}$ serve as a notable example where our results can be applied very effectively.
Theorem 13.3.
Let $m\in\mathbb{N}$. Assume A.1 for $m$ and A.2. Then, $(\mathscr{E}^{\mu^{[m]}},\overline{\mathscr{D}}^{\mu^{[m]}}_{\circ})$ is a quasi-regular Dirichlet form on $L^{2}(\mu^{[m]})$.
Proof.
Our goal is to prove ((QR1)), ((QR2)), and ((QR3)) for $(\mathscr{E}^{\mu^{[m]}},\overline{\mathscr{D}}^{\mu^{[m]}}_{\circ})$ on $L^{2}(\mu^{[m]})$.
Assumptions ((QR2)) and ((QR3)) are obvious. Indeed, $(\mathscr{E}^{\mu^{[m]}},\overline{\mathscr{D}}^{\mu^{[m]}}_{\circ})$ is the closure of $(\mathscr{E}^{\mu^{[m]}},\mathscr{D}_{\circ}^{\mu^{[m]}})$ on $L^{2}(\mu^{[m]})$ and $\mathscr{D}_{\circ}^{\mu^{[m]}}\subset C((\mathbb{R}^{d})^{m}\times\mathsf{S})$. Each element of $\mathscr{D}_{\circ}^{\mu^{[m]}}$ is continuous, which yields ((QR2)). From $\mathscr{D}_{\circ}^{\mu^{[m]}}$, we can extract a countable dense subset, which yields ((QR3)). Thus, it only remains to prove ((QR1)).
From Theorem 13.2, ((QR1)) holds for $\mu$. Hence, there exists an $\mathscr{E}^{\mu}$-nest $\{\mathsf{K}_{\mathsf{r}}\}_{\mathsf{r}\in\mathbb{N}}$ consisting of compact sets. Then, using the characterization of compact sets in $\mathsf{S}$, we have an increasing sequence of increasing sequences of natural numbers $a_{\mathsf{r}}=\{a_{\mathsf{r}}(Q)\}_{Q\in\mathbb{N}}$, $\mathsf{r}\in\mathbb{N}$, such that
$\mathsf{K}_{\mathsf{r}}\subset\mathsf{K}[a_{\mathsf{r}}],$
where $\mathsf{K}[a_{\mathsf{r}}]$ be the set defined by (2.44). Then, clearly $\{\mathsf{K}[a_{\mathsf{r}}]\}$ is a compact $\mathscr{E}^{\mu}$-nest. Let $a_{\mathsf{r}}^{m}=\{a_{\mathsf{r}}(Q)-m\}_{Q\in\mathbb{N}}$ and
$\mathsf{K}_{\mathsf{r}}^{[m]}=\overline{S}_{\mathsf{r}}^{m}\times\mathsf{K}[a_{\mathsf{r}}^{m}].$
Then, $\mathsf{K}_{\mathsf{r}}^{[m]}$, $\mathsf{r}\in\mathbb{N}$, is an increasing sequence of compact sets in $(\mathbb{R}^{d})^{m}\times\mathsf{S}$. Let $\mu_{\mathsf{r}}=\mu(\cdot\cap\mathsf{K}[a_{\mathsf{r}}])$. Let $\mu^{[m]}_{\mathsf{r}}$ be the $m$-reduced Campbell measure of $\mu_{\mathsf{r}}$. Then, the $m$-reduced Campbell measure $\mu^{[m]}$ is defined as the increasing limit of $\mu^{[m]}_{\mathsf{r}}$ as $\mathsf{r}\to\infty$ and
$\mu^{[m]}_{\mathsf{r}}=\mu^{[m]}(\cdot\cap\mathsf{K}_{\mathsf{r}}^{[m]}).$
We want to prove $\{\mathsf{K}_{\mathsf{r}}^{[m]}\}$ is an $\mathscr{E}^{\mu^{[m]}}$-nest.
Let $\tilde{\mu}_{\mathsf{r}}^{m}=\sum_{n=m}^{\infty}n(n-1)\cdots(n-m+1)\mu(\cdot\cap\mathsf{S}_{\mathsf{r}}^{n}\cap\mathsf{K}[a_{\mathsf{r}}])$. The number of particles in $S_{\mathsf{r}}$ under $\mu^{[m]}(\cdot\cap\mathsf{K}_{\mathsf{r}}^{[m]})$ is bounded by $a_{\mathsf{r}}(\mathsf{r})$ by $a_{\mathsf{r}}=\{a_{\mathsf{r}}(Q)\}_{Q\in\mathbb{N}}$. Hence,
$\tilde{\mu}_{\mathsf{r}}^{m}=\sum_{n=m}^{a_{\mathsf{r}}(\mathsf{r})}n(n-1)\cdots(n-m+1)\mu(\cdot\cap\mathsf{S}_{\mathsf{r}}^{n}\cap\mathsf{K}[a_{\mathsf{r}}]).$ (13.11)
Let
$\mathscr{E}^{\tilde{\mu}_{\mathsf{r}}^{m}}(f,g)=\int_{\mathsf{S}}\mathbb{D}_{\mathfrak{a}}[f,g]d\tilde{\mu}_{\mathsf{r}}^{m},$
$\overline{\mathscr{D}}_{\circ}^{\tilde{\mu}_{\mathsf{r}}^{m}}=\{f\in\overline{\mathscr{D}}_{\circ}^{\mu};\mathscr{E}^{\tilde{\mu}_{\mathsf{r}}^{m}}(f,f)<\infty,\,f\in L^{2}(\tilde{\mu}_{\mathsf{r}}^{m})\}.$ (13.12)
For a closed set $A$, we set
$\overline{\mathscr{D}}_{\circ}^{\tilde{\mu}_{\mathsf{r}}^{m}}(A)=\{f\in\overline{\mathscr{D}}_{\circ}^{\tilde{\mu}_{\mathsf{r}}^{m}};f(\mathsf{s})=0\text{ for $\tilde{\mu}_{\mathsf{r}}^{m}$-a.e.\,}\mathsf{s}\in A^{c}\},$
$\overline{\mathscr{D}}_{\circ}^{\mu^{[m]}_{\mathsf{r}}}(A)=\{f\in\overline{\mathscr{D}}_{\circ}^{\mu^{[m]}_{\mathsf{r}}};f(\mathsf{s})=0\text{ for $\mu^{[m]}_{\mathsf{r}}$-a.e.\,}\mathsf{s}\in A^{c}\}.$
Let
$\overline{\mathscr{D}}_{\circ}^{\tilde{\mu}_{\mathsf{r}}^{m}}(\{\mathsf{K}[a_{\mathsf{r}}]\})=\bigcup_{\mathsf{r}=1}^{\infty}\overline{\mathscr{D}}_{\circ}^{\tilde{\mu}_{\mathsf{r}}^{m}}(\mathsf{K}[a_{\mathsf{r}}]),$
$\overline{\mathscr{D}}_{\circ}^{\mu^{[m]}_{\mathsf{r}}}(\{\overline{S}_{\mathsf{r}}^{m}\times\mathsf{K}[a_{\mathsf{r}}^{m}]\})=\bigcup_{\mathsf{r}=1}^{\infty}\overline{\mathscr{D}}_{\circ}^{\mu^{[m]}_{\mathsf{r}}}(\overline{S}_{\mathsf{r}}^{m}\times\mathsf{K}[a_{\mathsf{r}}^{m}]).$
From (13.11) and (13.12), $\overline{\mathscr{D}}_{\circ}^{\tilde{\mu}_{\mathsf{r}}^{m}}(\{\mathsf{K}[a_{\mathsf{r}}]\})$ is dense in $\overline{\mathscr{D}}_{\circ}^{\tilde{\mu}_{\mathsf{r}}^{m}}$ with respect to the norm
$\lVert f\rVert_{\mathscr{E}_{1}^{\tilde{\mu}_{\mathsf{r}}^{m}}}=\{\mathscr{E}^{\tilde{\mu}_{\mathsf{r}}^{m}}(f,f)+\int_{\mathsf{S}}\lvert f\rvert^{2}d\tilde{\mu}_{\mathsf{r}}^{m}\}^{1/2}.$
Let $f\otimes g\in\mathscr{D}_{\circ}^{\mu^{[m]}}$ and $f\otimes h\in C_{0}^{\infty}((\mathbb{R}^{d})^{m})\otimes\overline{\mathscr{D}}_{\circ}^{\tilde{\mu}_{\mathsf{r}}^{m}}(\mathsf{K}[a_{\mathsf{r}}^{m}])$. Then, $f\otimes h\in\mathscr{D}_{\circ}^{\mu^{[m]}}$.
Let $\mathsf{r}\in\mathbb{N}$ such that f(x)=0f({x})=0 for ${x}\notin S_{\mathsf{r}}^{m}$ and that $g$ is $\sigma[\pi_{S_{\mathsf{r}}}]$-measurable. Let $(f\otimes g)_{\mathrm{sym}}$ be the symmetrization of $f\otimes g$ defined by $(f\otimes g)_{\mathrm{sym}}(\mathsf{s})=\sum_{\sigma\in\mathfrak{S}_{m}}f(\sigma({x}))g(\mathsf{t})$, where $\mathsf{s}=\sum_{i=1}^{m}\delta_{{x}_{i}}+\mathsf{t}$ and $\mathfrak{S}_{m}$ is the symmetric group of order $m$. Then, $h_{\mathrm{sym}}\in\mathsf{K}[a_{\mathsf{r}}]$.
Let $c_{13.1}$ and $c_{13.2}$ be constants defined by
[…]
By a straightforward calculation, we have, writing $\mathscr{E}^{\mu^{[m]}}_{1}(f,f)=\mathscr{E}^{\mu^{[m]}}_{1}(f)$,
$\mathscr{E}^{\mu^{[m]}}_{1}(f\otimes g-f\otimes h)=$ $\mathscr{E}^{\mu^{[m]}}_{1}(f\otimes(g-h))$
[…]
Let $\mathscr{D}^{[m]}=\cup_{\mathsf{r}}C_{0}^{\infty}(S_{\mathsf{r}}^{m})\otimes\mathscr{D}_{\circ}(\mathsf{K}[a_{\mathsf{r}}^{m}])$. Recall that $a_{\mathsf{r}}^{m}=\{a_{\mathsf{r}}(Q)-m\}_{Q\in\mathbb{N}}$. Note that $\mathscr{D}_{\circ}(\{\mathsf{K}[a_{\mathsf{r}}]\})$ is dense in $\overline{\mathscr{D}}_{\circ}^{\mu}$ because $\{\mathsf{K}[a_{\mathsf{r}}]\}$ is an $\mathscr{E}^{\mu}$-nest. Then, $\mathscr{D}^{[m]}$ is dense in $\mathscr{D}_{\circ}^{\mu^{[m]}}$ from (13.13). Hence, $\mathscr{D}^{[m]}$ is dense in $\overline{\mathscr{D}}^{\mu^{[m]}}_{\circ}$, which implies ((QR1)). ∎
13.4. Weak solution and the IFC condition to ISDEs.
In Subsection 13.4, we quote the IFC condition from [34]. Let
$\mathsf{S}_{\mathrm{s,i}}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\mathbb{R}^{d})=\infty,\mathsf{s}(\{s\})\in\{0,1\}\text{ for all }s\in\mathbb{R}^{d}\}.$

### v4

@@ A Dirichlet form $(\mathscr{E},\mathscr{D})$ on $L^{2}(\nu)$ is quasi-regular if:
13.3. Representation of functions on $\mathsf{S}$ and correlation functions
Definition 13.5.
A symmetric, locally integrable function $\rho_{\nu}^{m}$ on $(\mathbb{R}^{d})^{m}$ is called the $m$-point correlation function of an RPF $\nu$ (with respect to Lebesgue measure) if
$\int_{A_{1}^{k_{1}}\times\cdots\times A_{n}^{k_{n}}}\rho_{\nu}^{m}({x}^{1},\ldots,{x}^{m})\,d{x}^{1}\cdots d{x}^{m}=\int_{\mathsf{S}}\prod_{i=1}^{n}\frac{\mathsf{s}(A_{i})!}{(\mathsf{s}(A_{i})-k_{i})!}\,\nu(d\mathsf{s}),$
for any disjoint bounded measurable sets $A_{1},\ldots,A_{n}$ and nonnegative integers $k_{1}+\cdots+k_{n}=m$. Here, $\mathsf{s}(A)$ denotes the number of particles in $A$ when $\mathsf{s}=\sum_{i}\delta_{s^{i}}$, and the fraction is interpreted as zero if $\mathsf{s}(A_{i})-k_{i}<0$.
Let $\mathbf{a}=\{a_{\mathsf{q}}\}_{\mathsf{q}\in\mathbb{N}}$ be a family of sequences $a_{\mathsf{q}}=\{a_{\mathsf{q}}(R)\}_{R\in\mathbb{N}}$ of natural numbers. Assume that, for all $\mathsf{q},R\in\mathbb{N}$,
$a_{\mathsf{q}}(R)<a_{\mathsf{q}}(R+1),\quad a_{\mathsf{q}}(R)<a_{\mathsf{q}+1}(R),\quad a_{\mathsf{q}}^{+}(R)<a_{\mathsf{q}+1}(R),$ (13.3)
where $a_{\mathsf{q}}^{+}(R)=1+a_{\mathsf{q}}(R+1)$. For $a_{\mathsf{q}}=\{a_{\mathsf{q}}(R)\}_{R\in\mathbb{N}}$, let
$\mathsf{K}[a_{\mathsf{q}}]$ $=\bigcap_{R=1}^{\infty}\{\mathsf{s}\in\mathsf{S};\mathsf{s}(S_{R})\leq a_{\mathsf{q}}(R)\}.$ (13.4)
Then $\mathsf{K}[a_{\mathsf{q}}]$ is a compact set in $\mathsf{S}$ for each $\mathsf{q}\in\mathbb{N}$ such that
$\mathsf{K}[a_{\mathsf{q}}]\subset\mathsf{K}[\mathit{a}_{\mathsf{q}}^{+}]\subset\mathsf{K}[a_{\mathsf{q}+1}].$ (13.5)
Let $\nu$ be an RPF and $m\in\mathbb{N}$. Let $\nu_{\mathsf{q}}=\nu(\cdot\cap\mathsf{K}[a_{\mathsf{q}}])$. Suppose that $\nu$ has density functions. Then by (13.4), $\nu_{\mathsf{q}}$ has the $m$-point correlation function and the $m$-reduced Campbell measure $\nu^{[m]}_{\mathsf{q}}$ similarly as (2.35). We define $\nu^{[m]}$ as the increasing limit of $\nu^{[m]}_{\mathsf{q}}$:
$\quad\nu^{[m]}(A)=\lim_{\mathsf{q}\to\infty}\nu^{[m]}_{\mathsf{q}}(A),\quad A\in\mathscr{B}((\mathbb{R}^{d})^{m}\times\mathsf{S}).$ (13.6)
Note that $\nu^{[m]}$ is independent of the particular choice of compact sets $\mathsf{K}[a_{\mathsf{q}}]$.
Let $\|h\|:=\|h\|_{L^{1}(S_{Q}^{m}\times\mathsf{S},\nu^{[m]})}+\|\nabla_{{x}}h\|_{L^{1}(S_{Q}^{m}\times\mathsf{S},\nu^{[m]})}$. We note that, in the definition of $\|\cdot\|$, the differential is taken only with respect to the $x$-variable, and not with respect to $\mathsf{s}\in\mathsf{S}$. This point is very different from the energy norms used in Theorems 13.1 and 13.2.
Lemma 13.6.
The set $\mathscr{D}_{\circ\mathrm{b}}^{[m]}$ is dense in $\mathscr{D}_{\bullet\mathrm{b}}^{[m]}$ with respect to $\|\cdot\|$, $Q\in\mathbb{N}$.
@@ 13.3. Representation of functions on $\mathsf{S}$ and correlation functions
Recall that $\mathscr{D}_{\circ\mathrm{b}}^{[m]}=C_{0}^{\infty}((\mathbb{R}^{d})^{m})\otimes\mathscr{D}_{\circ\mathrm{b}}$ and $\mathscr{D}_{\bullet\mathrm{b}}^{[m]}=C_{0}^{\infty}((\mathbb{R}^{d})^{m})\otimes\mathscr{D}_{\bullet\mathrm{b}}$.
From the proof of Lemma 2.4 in [24], we easily see that both $\mathscr{D}_{\circ\mathrm{b}}$ and $\mathscr{D}_{\bullet\mathrm{b}}$ are dense in $L^{2}(\nu)$. Hence, we can prove Lemma 13.6 from this. ∎
For $R\in\mathbb{N}$ and $m\in\{0\}\cup\mathbb{N}$, set $\mathsf{S}_{R}^{m}=\{\mathsf{s}\in\mathsf{S};\mathsf{s}(S_{R})=m\}$. For $A\in\mathscr{B}(\mathsf{S})$, define $\pi_{A}(\mathsf{s})=\mathsf{s}(\cdot\cap A)$. Let $S_{R}^{m}$ denote the $m$-fold product of $S_{R}$.
Definition 13.6.
i For $\mathsf{s}\in\mathsf{S}$, a coordinate of $\mathsf{s}$ on $S_{R}$ is a tuple $\mathbf{x}_{R}(\mathsf{s})=(x_{R}^{i}(\mathsf{s}))_{i}\in\sqcup_{m=1}^{\infty}S_{R}^{m}$ such that $\pi_{S_{R}}(\mathsf{s})=\sum_{i}\delta_{x_{R}^{i}(\mathsf{s})}$.
ii The restriction $\mathbf{x}_{R}^{m}(\mathsf{s})$ of $\mathbf{x}_{R}(\mathsf{s})$ on $S_{R}^{m}$ is called the $S_{R}^{m}$-coordinate.
iii Let $f:\mathsf{S}\to\mathbb{R}$ and $R,m\in\mathbb{N}$. We say that $f_{R,\mathsf{s}}^{m}$ is the $S_{R}^{m}$-representation of $f$ if $f_{R,\mathsf{s}}^{m}(\mathbf{x})$ is a function on $S_{R}^{m}$ such that
$f_{R,\cdot}^{m}:\mathsf{S}\times S_{R}^{m}\to\mathbb{R},\qquad(\mathsf{s},\mathbf{x})\mapsto f_{R,\mathsf{s}}^{m}(\mathbf{x}),$
and the following conditions hold:
$f_{R,\mathsf{s}}^{m}(\mathbf{x})\ \text{is symmetric in $\mathbf{x}$ for each $\mathsf{s}\in\mathsf{S}_{R}^{m}$},$
$f_{R,\mathsf{s}(1)}^{m}(\mathbf{x})=f_{R,\mathsf{s}(2)}^{m}(\mathbf{x})\ \text{if }\pi_{S_{R}^{c}}(\mathsf{s}(1))=\pi_{S_{R}^{c}}(\mathsf{s}(2)),$
$f_{R,\mathsf{s}}^{m}(\mathbf{x}_{R}(\mathsf{s}))=f(\mathsf{s})\ \text{for }\mathsf{s}\in\mathsf{S}_{R}^{m},$
$f_{R,\mathsf{s}}^{m}(\mathbf{x})=0\ \text{for }\mathsf{s}\notin\mathsf{S}_{R}^{m}.$
We define $f_{R,\mathsf{s}}$ on $\sqcup_{m\in\mathbb{N}}S_{R}^{m}$ by $f_{R,\mathsf{s}}(\mathbf{x})=f_{R,\mathsf{s}}^{m}(\mathbf{x})$ for $\mathbf{x}\in S_{R}^{m}$.
$f(\mathsf{s})=f_{R,\mathsf{s}}(\mathbf{x}_{R}(\mathsf{s})).$
By convention, $\mathbf{x}_{R}(\mathsf{s})=\emptyset$ for $\mathsf{s}\in\mathsf{S}_{R}^{0}$, and $f_{R,\mathsf{s}}$ is constant on $\mathsf{S}_{R}^{0}$.
For a bounded set $A$, one defines $\mathbf{x}_{A}^{m}(\mathsf{s})$, $f_{A}^{m}(\mathbf{x})$ and $f_{A,\mathsf{s}}(\mathbf{x})$ analogously by replacing $S_{R}$ with $A$.
13.4. Quasi-regularity of $m$-labeled Dirichlet forms
Let the coefficient $\mathfrak{a}$ satisfy (2.31) and $\mathbb{D}_{\mathfrak{a}}^{[m]}$ be as (7.1). Let
$\mathscr{E}^{[m]}(f,g)=\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}\mathbb{D}_{\mathfrak{a}}^{[m]}[f,g]d\nu^{[m]},$
$\mathscr{D}_{\circ}^{[m]}=\{f\in C_{0}^{\infty}((\mathbb{R}^{d})^{m})\otimes\mathscr{D}_{\circ};\mathscr{E}^{[m]}(f,f)<\infty,\,f\in L^{2}(\nu^{[m]})\}.$
Although $\mathscr{E}^{[m]}$, $\mathscr{D}_{R\bullet}^{[m]}$, and $\mathscr{D}_{\circ}^{[m]}$ depend on $\mathfrak{a}$ and $\mu$, we suppress this dependence in the notation. We make assumptions:
A.1 $(\mathscr{E}^{[m]},\mathscr{D}_{\circ}^{[m]})$ is closable on $L^{2}(\nu^{[m]})$.
A.2 $\int_{\mathsf{S}}\mathsf{s}(S_{R})\nu(d\mathsf{s})<\infty$ for each $R\in\mathbb{N}$.
A.3 $\nu$ has a bounded $k$-density function on $\overline{S}_{R}^{k}$ for each $R,k\in\mathbb{N}$.
Let $(\mathscr{E}^{[m]},\overline{\mathscr{D}_{\circ}^{[m]}})$ be the closure of $(\mathscr{E}^{[m]},\mathscr{D}_{\circ}^{[m]})$ on $L^{2}(\nu^{[m]})$. We write $(\mathscr{E}^{\nu},\overline{\mathscr{D}}_{\circ}^{\nu})$ for $m=0$ since $\nu^{[0]}=\nu$.
Theorem 13.1 ([24, Th. 1]).
Assume A.1 for $m=0$, A.2, and A.3. Then the Dirichlet form $(\mathscr{E}^{\nu},\overline{\mathscr{D}}_{\circ}^{\nu})$ on $L^{2}(\nu)$ is quasi-regular.
Proof.
The theorem was proved in [24, Th. 1] under the assumption that the diffusion matrix $\mathfrak{a}$ is the identity. In the present setting, $\mathfrak{a}$ is uniformly elliptic and bounded by (2.31), and hence the same argument applies verbatim. ∎
Lemma 13.7.
Assume A.2. Then $\mathscr{D}_{\circ}^{\nu}$ is dense in $L^{2}(\nu)$.
Proof.
The assertion was proved in the proof of Lemma 2.5 in [17]. See the last line of the proof of Lemma 2.5 in [17]. ∎
In Theorem 13.2, we significantly relax the assumptions introduced in Theorem 13.1. Specifically, if $(\mathscr{E}^{[m]},\mathscr{D}_{\circ}^{[m]})$ is found to be closable on $L^{2}(\nu^{[m]})$, then the existence of a density function is no longer necessary.
@@ 13.4. Quasi-regularity of $m$-labeled Dirichlet forms
Assume A.1 and A.2. Then $(\mathscr{E}^{[m]},\overline{\mathscr{D}_{\circ}^{[m]}})$ is a quasi-regular Dirichlet form on $L^{2}(\nu^{[m]})$.
@@
Assumptions ((Q1)) and ((Q3)) follow from A.1 and A.2 by exactly the same arguments as those used in the proofs of [24, Th. 1] and [26, Lem. 2.3]. Thus, it only remains to prove ((Q2)).
For $m=0$, A.3 is used only in the proof of [24, Lem. 2.4] to establish Lemma 13.7. Hence, ((Q2)) holds for $m=0$.
For $m\in\mathbb{N}$, the result follows from [26, Lem. 2.3] by reducing to the case $m=0$, and we therefore omit the proof. ∎
13.5. Weak solution and the IFC condition to ISDEs
In Subsection 13.5, we quote the IFC condition from [34].